OpenStudy (linn99123):

NEED HELP FAST MEDAL AND FAN what is the solution to the proportion x/45=8/18

OpenStudy (anonymous):

x/45=8/18 Simple and best practice solution for x/45=8/18 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for x/45=8/18 equation: x in (-oo:+oo) x/45 = 8/18 // - 8/18 x/45-(8/18) = 0 x/45-4/9 = 0 1/45*x-4/9 = 0 // + 4/9 1/45*x = 4/9 // : 1/45 x = 4/9/1/45 x = 20 x = 20

undeadknight26 (undeadknight26):

think this is 20/45 because" 45/18 = 2.5 8 x 2.4 = 20.

undeadknight26 (undeadknight26):

o.o or that...

OpenStudy (linn99123):

LOL knight xD

undeadknight26 (undeadknight26):

no wait i made a typo... 8 x 2.5 = 20... But im sure that sashmay did it the right way XD

OpenStudy (anonymous):

haha thanks. Im in Calculus